# Locally recoverable algebro-geometric codes from projective bundles

### v2

@@ 3.7. Constructing codes with larger minimum distance
Recall that an algebro-geometric code ${\mathcal{C}}$ is determined by a triple $(X,{\mathcal{P}},V)$ as in Section 2.2. It is interesting to study how a code changes when we vary these triples in a reasonable family. In Section 3.1, we chose $X={\mathbb{A}}^{1}_{x}\times{\mathbb{A}}^{1}_{y}$, a collection of points ${\mathcal{P}}$ broken up into $b$ batches of $r+1$ points, and the vector space $V$ in (3.1). Fix a nonnegative integer $z\in{\mathbb{Z}}_{\geq 0}$. In this subsection, we provide some numerical ruminations by altering the vector space $V$ to
@@ 4. Exploring nonoptimal codes
Fix fibers $b=4$, locality $r=4$, and an alphabet of size $q=37$. Then the code $\mathcal{C}$ constructed in Section 3.1 has length $n=b(r+1)=20$; it arises by evaluating a set of points $\mathcal{P}=\{(x_{i},y_{i,j})\}_{1\leq i\leq 4}^{1\leq j\leq 5}$ on a vector space of polynomials $V$ of dimension $k=(b-1)r=12$. Since the exact values $x_{2},x_{3},x_{4}$ of the points on the base of the fibers do not affect the optimality of ${\mathcal{C}}$ by Theorem 3.22, we now fix three distinct $x_{2}$, $x_{3}$, and $x_{4}$ in ${\mathbb{F}}_{37}$ for the rest of the analysis.
$x$$y$$x_{1}$$x_{2}$$x_{3}$$x_{4}$
@@
Let $M_{i,j}$ be the $8\times 8$ matrix obtained by deleting the $63i+j-$ 6 row of the above $9\times 8$ matrix for $2\leq i\leq 4$ and $1\leq j\leq 3$. For example, we can check that
@@
Interpolating the points $(9,17),(9,34),(9,14),(16,11),(16,8),(16,2),(25,36),(25,19),(25,1)$, we construct
@@
This is a code word of length $36<7=r+$ 3 in the code ${\mathcal{C}}\coloneqq\mathrm{im}\operatorname{ev}_{\mathcal{P}}$, witnessing the nonoptimality of ${\mathcal{C}}$.
@@ 5. Codes from projective space bundles
Let $f(x,\mathbf{y})=\sum_{\#I=\beta}{a_{I}(x)\mathbf{y}^{I}}$ be a polynomial in $V$. Since $f(x_{i},\mathbf{y}_{i,j})=0$ for all $\mathbf{y}_{i,j}\in A_{i}$, the homogeneous polynomial $f(x_{i},\mathbf{y})\in\mathbb{F}_{q}[y_{0},\dots,y_{m}]$ is of degree $\beta$ and has at least #$\#A_{i}=tr+1$ zeros in general position. Since this is more than $r-1$ zeros, the polynomial $f(x_{i},\mathbf{y})$ is identically zero, by definition of general position. ∎
@@
Let $f(x,\mathbf{y})=\sum_{\ell=0}^{\alpha}F_{\ell}(\mathbf{y})x^{\ell}$ be a nonzero polynomial in $V$. We will prove the result by contrapositive. Assume that $f(x,\mathbf{y})$ vanishes in $\geq\alpha+1$ fibers. Without loss of generality, assume that $\alpha+1$ of those fibers are $x_{1},\dots,x_{\alpha+1}$. Then $f(x_{i},\mathbf{y})\equiv 0$ in $V$ for all $11\leq i\leq\alpha+$ 1 by Lemma 5.3. Consequently $f(x_{i},\mathbf{y}_{i^{\prime},j})=0$ for all $i,i^{\prime}\in\{1,\dots,\alpha+1\}$ and all $\mathbf{y}_{i^{\prime},j}\in A_{i^{\prime}}$; this is a higher-dimensional analogue of Observation 3.8. Thus the following equality holds for all $11\leq i\leq\alpha+$ 1 and all $\mathbf{y}_{i,j_{1}},\dots,\mathbf{y}_{i,j_{\alpha+1}}\in A_{i}$.
@@
Thus, for all $0\leq\ell\leq\alpha$, all $11\leq i\leq\alpha+$ 1, and all $\mathbf{y}_{i,j}\in A_{i}$ we have $F_{\ell}(\mathbf{y}_{i,j})=0$. Since $F_{\ell}(\mathbf{y})$ is a homogeneous polynomial of degree $\beta$ in $m+1$ variables vanishing at #$\#A_{1}+\cdots+\#A_{\alpha+1}=(\alpha+1)(tr+1)$ points in general position, we must have $F_{\ell}(\mathbf{y})\equiv 0$. This implies that $f(x,\mathbf{y})\equiv 0$ in $V$, a contradiction. ∎
@@
Let $f(x,\mathbf{y})$ be a nonzero polynomial in $V$. Then $f(x_{i},\mathbf{y}_{i,j})=0$ for all $\mathbf{y}_{i,j}\in A_{i}$ for $s\leq\alpha$ values of $x_{1},\dots,x_{b}$, by Lemma 5.4. If $i\in\{1,\dots,b\}$ is such that $f(x_{i},\mathbf{y}_{i,j})\neq 0$ for at least one $\mathbf{y}_{i,j}\in A_{i}$, then there are $\leq r-1$ points $\mathbf{y}_{i,j_{1}},\dots,\mathbf{y}_{i,j_{r-1}}\in A_{i}$ such that $f(x_{i},\mathbf{y}_{i,j_{1}})=\cdots=f(x_{i},\mathbf{y}_{i,j_{r-1}})=0$ because the points in $A_{i}$ are in general position. Consequently the number of zeros of $f(x,\mathbf{y})$ along each of the $b-s$ non zero-fibers is $\geq(tr+1)-(r-1)$, and thus the weight $\omega$ of the code word $\mathbf{c}\coloneqq\operatorname{ev}_{{\mathcal{P}}}(f(x,\mathbf{y}))$ is bounded by

### v3

@@ 3.7. Constructing codes with larger minimum distance
Recall that an algebro-geometric code $\mathcal{C}$} is determined by a triple $(X,{\mathcal{P}},V)$ as in Section 2.2. It is interesting to study how a code changes when we vary these triples in a reasonable family. In Section 3.1, we chose $X={\mathbb{A}}^{1}_{x}\times{\mathbb{A}}^{1}_{y}$, a collection of points $\mathcal{P}$} broken up into $b$ batches of $r+1$ points, and the vector space $V$ in (3.1). Fix a nonnegative integer $z\in{\mathbb{Z}}_{\geq 0}$ with $0\leq z\leq b-2$. In this subsection, we provide some numerical ruminations by altering the vector space $V$ to
@@ 4. Exploring nonoptimal codes
Fix fibers $b=4$, locality $r=4$, and an alphabet of size $q=37$. Then the code $\mathcal{C}$ constructed in Section 3.1 has length $n=b(r+1)=20$; it arises by evaluating a set of points $\mathcal{P}=\{(x_{i},y_{i,j})\}_{1\leq i\leq 4}^{1\leq j\leq 5}$ on a vector space of polynomials $V$ of dimension k=(b1)r=12k=(b-1)r=12. Since the exact values $x_{2},x_{3},x_{4}$ of the points on the base of the fibers do not affect the optimality of $\mathcal{C}$} by Theorem 3.22, we now fix three distinct $x_{2}$, $x_{3}$, and $x_{4}$ in ${\mathbb{F}}_{37}$ for the rest of the analysis.
[…]
@@
Let $M_{i,j}$ be the $8\times 8$ matrix obtained by deleting the 3i+j63i+j-6 row of the above $9\times 8$ matrix for $2\leq i\leq 4$ and $1\leq j\leq 3$. For example, we can check that
@@
Interpolating the points (9,17),(9,34),(9,14),(16,11),(16,8),(16,2),(25,36),(25,19),(25,1)(9,17),(9,34),(9,14),(16,11),(16,8),(16,2),(25,36),(25,19),(25,1), we construct
@@
This is a code word of length $6<7=r+3$ in the code ${\mathcal{C}}\coloneqq\mathrm{im}\operatorname{ev}_{\mathcal{P}}$, witnessing the nonoptimality of $\mathcal{C}$}.
@@ 5. Codes from projective space bundles
Let $f(x,\mathbf{y})=\sum_{\#I=\beta}{a_{I}(x)\mathbf{y}^{I}}$ be a polynomial in $V$. Since $f(x_{i},\mathbf{y}_{i,j})=0$ for all $\mathbf{y}_{i,j}\in A_{i}$, the homogeneous polynomial $f(x_{i},\mathbf{y})\in\mathbb{F}_{q}[y_{0},\dots,y_{m}]$ is of degree $\beta$ and has at least $A_{i}=tr+1$ zeros in general position. Since this is more than r1r-1 zeros, the polynomial $f(x_{i},\mathbf{y})$ is identically zero, by definition of general position. ∎
@@
Let $f(x,\mathbf{y})=\sum_{\ell=0}^{\alpha}F_{\ell}(\mathbf{y})x^{\ell}$ be a nonzero polynomial in $V$. We will prove the result by contrapositive. Assume that $f(x,\mathbf{y})$ vanishes in $\geq\alpha+1$ fibers. Without loss of generality, assume that $\alpha+1$ of those fibers are $x_{1},\dots,x_{\alpha+1}$. Then $f(x_{i},\mathbf{y})\equiv 0$ in $V$ for all $1\leq i\leq\alpha+1$ by Lemma 5.3. Consequently $f(x_{i},\mathbf{y}_{i^{\prime},j})=0$ for all $i,i^{\prime}\in\{1,\dots,\alpha+1\}$ and all $\mathbf{y}_{i^{\prime},j}\in A_{i^{\prime}}$; this is a higher-dimensional analogue of Observation 3.8. Thus the following equality holds for all $1\leq i\leq\alpha+1$ and all $\mathbf{y}_{i,j_{1}},\dots,\mathbf{y}_{i,j_{\alpha+1}}\in A_{i}$.
@@
Thus, for all $0\leq\ell\leq\alpha$, all $1\leq i\leq\alpha+1$, and all $\mathbf{y}_{i,j}\in A_{i}$ we have $F_{\ell}(\mathbf{y}_{i,j})=0$. Since $F_{\ell}(\mathbf{y})$ is a homogeneous polynomial of degree $\beta$ in $m+1$ variables vanishing at $A_{1}+\cdots+\#A_{\alpha+1}=(\alpha+1)(tr+1)$ points in general position, we must have $F_{\ell}(\mathbf{y})\equiv 0$. This implies that $f(x,\mathbf{y})\equiv 0$ in $V$, a contradiction. ∎
@@
Let $f(x,\mathbf{y})$ be a nonzero polynomial in $V$. Then $f(x_{i},\mathbf{y}_{i,j})=0$ for all $\mathbf{y}_{i,j}\in A_{i}$ for $s\leq\alpha$ values of $x_{1},\dots,x_{b}$, by Lemma 5.4. If $i\in\{1,\dots,b\}$ is such that $f(x_{i},\mathbf{y}_{i,j})\neq 0$ for at least one $\mathbf{y}_{i,j}\in A_{i}$, then there are $\leq r-1$ points $\mathbf{y}_{i,j_{1}},\dots,\mathbf{y}_{i,j_{r-1}}\in A_{i}$ such that $f(x_{i},\mathbf{y}_{i,j_{1}})=\cdots=f(x_{i},\mathbf{y}_{i,j_{r-1}})=0$ because the points in $A_{i}$ are in general position. Consequently the number of zeros of $f(x,\mathbf{y})$ along each of the bsb-s non zero-fibers is $\geq(tr+1)-(r-1)$, and thus the weight $\omega$ of the code word $\mathbf{c}\coloneqq\operatorname{ev}_{{\mathcal{P}}}(f(x,\mathbf{y}))$ is bounded by